A random variable is normally distributed with a mean of u=50 and a standard deviation of a=5 .The probability that the random variable will assume a value greater than 40 is
1. 0.9772
2. 0.0228
3. 0.4207
4. 0.5793
5. 0.4920
1
Expert's answer
2020-11-09T19:38:37-0500

"P(X>40)=P(Z>\frac{40-50}{5})"

"=P(Z>-2)=1-P(Z<-2)"

"=1-0.0228=0.9772"

Choice 1 is the correct answer.